What is the probability that a random 2 digit number is a multiple of 7?
Found 2 solutions by stanbon, robertb:
Answer by stanbon(75887)

What is the probability that a random 2 digit number is a multiple of 7?
----
# of 2-digit numbers: 99-10+1 = 90
----
# that are a multiple of 7
7*2,7*3,...,7*12
11 such numbers
---
P(a random 2-digit number is a multiple of 7) = 11/90
===================
Cheers,
Stan H.
===========

Answer by robertb(5830)

Without loss of generality, assume only positive integers. There are 90 2-digit positive integers, while there are 13 2-digit positive integers divisible by 7.
Hence the probability is 13/90.
Why 13 2-digit positive integers divisible by 7?
98 = 14 + (n-1)7, from the arithmetic sequence formula.
==> 84 = (n-1)7
==> 12 = n-1
==> n = 13.